# Kiloparsec-scale turbulence driven by reionization may grow intergalactic magnetic fields

### v1

@@ D.1.1 Spatial resolution
D.1.2 Box size & local density
Our fiducial simulations are run in extremely small boxes relative to typical cosmological simulations, sampling a patch of the IGM at the cosmic mean density. To study the effect of box size on our results, we ran an $L=0.5$ $h^{-1}$pc, flash-ionized box with $N=2048^{3}$, matching our fiducial cell size of $0.244$ $h^{-1}$kpc. We show a slice through the temperature field in this run at $z=5$ in Figure 8, alongside the same for the two smaller boxes shown in Figure 7, re-sized to scale. The upper-right quadrant of the $0.5$ $h^{-1}$pc slice, which we highlight, is mildly over-dense (a factor of $1.5$ relative to the rest of the slice). This sub-volume is dominated by a large filamentary structure that was recently destroyed by reionization.
Visually, there is no clear indication that the turbulent features are qualitatively different in boxes with different sizes. They appear to form at the same scale as in our fiducial box, and to fill roughly the same fraction of the slice. One might expect the destruction of more massive filaments to drive stronger turbulence due to the larger gas pressures driving the expansion. However, comparing the boxed region on the left to the right panels, there is no clear indication that this is the case. This suggests that our fiducial box size is sufficiently large to capture the essential properties of the turbulence.
Figure 8: Visualization of the effect of box size on IGM turbulence. On the left, we show a slice through the temperature in our $0.5$ $h^{-1}$pc, flash-ionized box at $z=5$. The upper-right quadrant, which is highlighted, is mildly over-dense relative to the rest of the slice. The small panels on the left are the same slices shown in Figure 7, re-sized to scale. We see no clear qualitative differences in turbulent features as a function of box-size or local over-density.
D.1.3 Sensitivity to modeling assumptions
@@
Figure 9 shows gas temperature slices at our fiducial resolution, each row varying a different simulation parameter. In the top row, we vary the redshift of reionization, $z_{\rm reion}$, and display the temperature map $\approx 300$ Myr after reionization (motivated by our picture that the turbulence is driven by how far sound waves can travel after ionization). The first panel shows our fiducial $z_{\rm reion}=7$, and the middle and right panels show $z_{\rm reion}=9$ and $5$, respectively. We see that turbulence is noticeably less (more) prominent in the $z_{\rm reion}=9$ ($5$) case. This is likely because denser, colder structures are able to form when the IGM re-ionizes at lower redshift, resulting in a more violent relaxation that drives turbulence more effectively. The gas at lower redshifts is also less dense and inverse Compton cools less quickly, such that the IGM takes more time to reach thermal and pressure equilibrium. Recent work supporting a late end to reionization [59, 60, 61, 62] suggests that as much as half of the IGM may have been neutral at $z=7$ and up to $20\%$ at $z=6$.
@@
Figure 9: Effect of different box scale physics on the emergence of IGM turbulence. Top row: effect of the reionization redshift. Higher (lower) $z_{\rm reion}$ results in less (more) turbulence, since more small-scale structure has formed in the neutral IGM at lower $z$. 2nd row: effect of X-ray pre-heating. From left to right, we impose pre-ionization minimum temperatures of $T_{\rm min}=10$, $100$, and $1000$ K at $z<15$ to crudely mimic the effects of pre-heating. These panels (and those below them) have $z_{\rm reion}=7$ and $z=5.5$. We see no difference with the unheated case for $T_{\rm min}=10$ K, but turbulence is greatly reduced for $T_{\rm min}=100$ K and eliminated entirely for $T_{\rm min}=1000$ K. 3rd row: effect of a small-scale cutoff in the dark matter power spectrum, parameterized by a WDM particle mass $m_{X}$ (see text). Turbulent features are missing entirely even for $m_{\rm X}=20$ keV. Bottom row: effect of the baryon-dark matter streaming velocity (see text). For $v_{\rm bc}^{\rm stream}=20$ km/s, there is little difference with the $v_{\rm bc}^{\rm stream}=0$ case, but higher $v_{\rm bc}^{\rm stream}$ leads to significantly lower levels of turbulence.
@@
In the second row of Figure 9, we show the effect of X-ray pre-heating of the pre-ionized IGM. We crudely model this by imposing a minimum temperature of $T_{\rm min}$ at $z\leq 15$ prior to reionization, following D’Aloisio et al. [5]. From left to right, we show $T_{\rm min}=10$, $100$, and $1000$ K for our fiducial $z_{\rm reion}=7$ at $z=5.5$. The left panel in the second row is almost indistinguishable from the fiducial case above it, suggesting that heating the IGM to $10$ K has almost no effect on turbulence as an unheated IGM. This is comparable to the current lower limits on the pre-heated IGM temperature from HERA at $z\approx 8$ [40]. However, for $T_{\rm min}=100$ K, turbulent features are considerably suppressed, although not entirely erased. This is consistent with our argument in §A.2 that pre-ionization gas temperatures much above $\approx 100$ K would prevent the seeding of turbulence. (One caveat is that our crude implementation of pre-heating – a temperature floor at $z<15$ may artificially increase its importance. Pre-heating likely starts later than this, such that for a given $T_{\rm min}$ at $z_{\rm reion}$, more small-scale structure forms and survives to $z_{\rm reion}$ than we find in our simulations.) For $T_{\rm min}=1000$ K, there is no sign of any turbulence.
@@
The last row shows the effect of the velocity offset between baryons and dark matter (streaming velocity, $v_{\rm bc}^{\rm stream}$) induced at recombination [68]. This effect suppresses the formation of structure on small scales by increasing the minimum DM halo mass required to accrete gas from the IGM. It has been shown to have a potentially significant effect on Pop III star formation and the 21 cm signal [69] and may affect the IGM even during reionization [70, 71]. We implement $v_{\rm bc}^{\rm stream}$ following the procedure used in [70]. We show results for $v_{\rm bc}^{\rm stream}=20$ km/s, $41$ km/s, and $65$ km/s at $z=1080$, the same values used in Cain et al. [70], which are $0.66$, $1.3$, and $\times2$ the RMS value of $\sigma_{\rm bc}=30$ km/s. Compared to the upper left-most panel, the case with $v_{\rm bc}^{\rm stream}=20$ km/s has only slightly reduced turbulence. However, at higher $v_{\rm bc}^{\rm stream}$, turbulent features become significantly less prominent, although they never disappear entirely. This raises the intriguing possibility that the strength of IGM magnetic fields may be a biased tracer of the large-scale $v_{\rm bc}^{\rm stream}$ field, which exhibits strong Baryon Acoustic Oscillation (BAO) features that are offset relative to the standard BAO used as a standard ruler [72]. If these magnetic fields then seed galactic magnetic fields, which then perhaps affect star formation, this could even result in a bias in the BAO peak position in galaxy clustering measurements.

### v2

@@ C.1.1 Spatial resolution
C.1.2 Box size & local density
Our fiducial simulations are run in extremely small boxes relative to typical cosmological simulations, sampling a patch of the IGM at the cosmic mean density. To study the effect of box size on our results, we ran an $L=0.5$ $h^{-1}$Mpc, flash-ionized box with $N=2048^{3}$, matching our fiducial cell size of $0.244$ $h^{-1}$kpc. We show a slice through the temperature field in this run at $z=5$ in Figure 6, alongside the same for the two smaller boxes shown in Figure 5, re-sized to scale. The upper-right quadrant of the $0.5$ $h^{-1}$Mpc slice, which we highlight, is mildly over-dense (a factor of $1.5$ relative to the rest of the slice). This sub-volume is dominated by a large filamentary structure that was recently destroyed by reionization.
Visually, there is no clear indication that the turbulent features are qualitatively different in boxes with different sizes. They appear to form at the same scale as in our fiducial box, and to fill roughly the same fraction of the slice. One might expect the destruction of more massive filaments to drive stronger turbulence due to the larger gas pressures driving the expansion. However, comparing the boxed region on the left to the right panels, there is no clear indication that this is the case. This suggests that our fiducial box size is sufficiently large to capture the essential properties of the turbulence. However, it is unclear whether much larger volumes, containing halos that would host star-forming galaxies, would have much different properties in over-densities hosting such halos.
Figure 6: Visualization of the effect of box size on IGM turbulence. On the left, we show a slice through the temperature in our $0.5$ $h^{-1}$Mpc, flash-ionized box at $z=5$. The upper-right quadrant, which is highlighted, is mildly over-dense relative to the rest of the slice. The small panels on the left are the same slices shown in Figure 5, re-sized to scale. We see no clear qualitative differences in turbulent features as a function of box-size or local over-density.
C.1.3 Vorticity
Figure 7: Visualization of the association of turbulence with vorticity. We show velocity (left) and vorticity (right) in $50$ $h^{-1}$kpc zoom-ins from our $T_{\rm min}=100$ K simulation with $\Delta x=122$ $h^{-1}$pc. The top row shows a turbulent region, and the bottom row a non-turbulent one.
It is well-known that for gas to exhibit turbulent flow, its vorticity (curl of the velocity) must be non-zero. Since the cosmological initial conditions are thought to have zero vorticity, the only way to source vorticity is if pressure and density gradients are not aligned, which occurs whenever pressure is not a unique function of density. This happens in our simulations after re-ionization as a direct results of the photo-heating and photoevaporation by cosmological ionization fronts [14, 17].
We show in Figure 7 that regions with high vorticity are, indeed, the sights of turbulence in our simulations. The top row shows a $50$ $h^{-1}$kpc zoom-in on a turbulent region in our $T_{\rm min}=100$ K simulation with $\Delta x=122$ $h^{-1}$pc. The left panel shows the magnitude of the gas velocity, and right shows the vorticity. The bottom row shows the same thing, but zoomed in on a region that is visibly not turbulent. We see that the turbulent case has high vorticity in regions where the gas flows are noticeably turbulent on the left. The gas velocities are also $\sim 20$ km/s here, near the sound speed. The non-turbulent region has somewhat slower velocity and small (but non-zero) vorticity. This confirms the expectation that turbulent regions should be associated with high-vorticity regions, and suggests that vorticity may be a useful metric for estimating the volume-filling factor of turbulence robustly, which we plan to study more carefully in future work.
C.1.4 Sensitivity to modeling assumptions
@@
Figure 8 shows gas temperature slices at our fiducial resolution, each row varying a different simulation parameter (at $z=5$, unless otherwise indicated). In the top row, we vary the redshift of reionization, $z_{\rm re}$, and display the temperature map $\approx 300$ Myr after reionization (motivated by our picture that the turbulence is driven by how far sound waves can travel after ionization). The first panel shows our fiducial $z_{\rm re}=7$, and the middle and right panels show $z_{\rm re}=9$ and $5$, respectively. We see that turbulence is noticeably less (more) prominent in the $z_{\rm re}=9$ (55) case. This is likely because denser, colder structures are able to form when the IGM re-ionizes at lower redshift, resulting in a more violent relaxation that drives turbulence more effectively. The gas at lower redshifts is also less dense and inverse Compton cools less quickly, such that the IGM takes more time to reach thermal and pressure equilibrium. Recent work supporting a late end to reionization [66, 67, 68, 69] suggests that as much as half of the IGM may have been neutral at $z=7$ and up to $20\%$ at $z=6$.
@@
Figure 8: Effect of different box scale physics on the emergence of IGM turbulence - all panels are shown at $z=5$, unless otherwise indicated. Top row: effect of the reionization redshift. Higher (lower) $z_{\rm re}$ results in less (more) turbulence, since more small-scale structure has formed in the neutral IGM at lower $z$. 2nd row: effect of X-ray pre-heating. From left to right, we impose pre-ionization minimum temperatures of $T_{\rm min}=10$, $100$, and $1000$ K at $z<15$ to crudely mimic the effects of pre-heating. We see no difference with the unheated case for $T_{\rm min}=10$ K, but turbulence is greatly reduced for $T_{\rm min}=100$ K and eliminated entirely for $T_{\rm min}=1000$ K. 3rd row: effect of a small-scale cutoff in the dark matter power spectrum, parameterized by a WDM particle mass $m_{X}$ (see text). Turbulent features are missing entirely even for $m_{\rm X}=20$ keV. Bottom row: effect of the baryon-dark matter streaming velocity (see text). For $v_{\rm bc}^{\rm stream}=20$ km/s, there is little difference with the $v_{\rm bc}^{\rm stream}=0$ case, but higher $v_{\rm bc}^{\rm stream}$ leads to significantly lower levels of turbulence.
@@
In the second row of Figure 8, we show the effect of X-ray pre-heating of the pre-ionized IGM. We crudely model this by imposing a minimum temperature of $T_{\rm min}$ at $z\leq 15$ prior to reionization, following D’Aloisio et al. [6]. From left to right, we show $T_{\rm min}=10$, $100$, and $1000$ K for our fiducial $z_{\rm re}=7$ at $z=5$. The left panel in the second row is almost indistinguishable from the fiducial case above it, suggesting that heating the IGM to $10$ K has almost no effect on turbulence as an unheated IGM. This is comparable to the current lower limits on the pre-heated IGM temperature from HERA at $z\approx 8$ [50]. However, for $T_{\rm min}=100$ K, turbulent features are considerably suppressed, although not entirely erased. This is consistent with our argument in §A.2 that pre-ionization gas temperatures much above $\approx 100$ K would prevent the seeding of turbulence. (One caveat is that our crude implementation of pre-heating – a temperature floor at $z<15$ - may artificially increase its importance. Pre-heating likely starts later than this, such that for a given $T_{\rm min}$ at $z_{\rm re}$, more small-scale structure forms and survives to $z_{\rm re}$ than we find in our simulations.) For $T_{\rm min}=1000$ K, there is no sign of any turbulence.
@@
The last row shows the effect of the velocity offset between baryons and dark matter (streaming velocity, $v_{\rm bc}^{\rm stream}$) induced at recombination [75]. This effect suppresses the formation of structure on small scales by increasing the minimum DM halo mass required to accrete gas from the IGM. It has been shown to have a potentially significant effect on Pop III star formation and the 21 cm signal [76] and may affect the IGM even during reionization [77, 78]. We implement $v_{\rm bc}^{\rm stream}$ following the procedure used in [77]. We show results for $v_{\rm bc}^{\rm stream}=20$ km/s, $41$ km/s, and $65$ km/s at $z=1080$, the same values used in Cain et al. [77], which are $0.66$, $1.3$, and $2.2\times$ the RMS value of $\sigma_{\rm bc}=30$ km/s. Compared to the upper left-most panel, the case with $v_{\rm bc}^{\rm stream}=20$ km/s has only slightly reduced turbulence. However, at higher $v_{\rm bc}^{\rm stream}$, turbulent features become significantly less prominent, although they never disappear entirely. This raises the intriguing possibility that the strength of IGM magnetic fields may be a biased tracer of the large-scale $v_{\rm bc}^{\rm stream}$ field, which exhibits strong Baryon Acoustic Oscillation (BAO) features that are offset relative to the standard BAO used as a standard ruler [79]. If these magnetic fields then seed galactic magnetic fields, which then perhaps affect star formation, this could even result in a bias in the BAO peak position in galaxy clustering measurements.
Next, we make these comparisons quantitative by estimating the number of eddy turnover times (Eq. 10) elapsed in each simulation by $z=4$. This is given by
$N_{\rm tau}(k,z=4)=\int_{z_{\rm re}}^{4}dz\frac{dt}{dz}\frac{1}{\tau_{\rm eddy}(k)}$ (13)
We show this quantity in Figure 9 for the same simulations shown in Figure 8. The top left panel compares simulations with $z_{\rm re}=5$, $7$, and $9$, and our high-resolution ($\Delta x=0.12$ $h^{-1}$) run with $z_{\rm re}=7$. We see that in the range of scales we previously identified as turbulent for our fiducial simulation ($300\lesssim k/[h{\rm Mpc}^{-1}]\lesssim 1000$), we find that $N_{\tau}$ scales approximately like $\propto k^{1}$, steeper than the $k^{2/3}$ scaling expected for the $E(k)\propto k^{-5/3}$ Kolmogorov scaling (see Eq. 10). The reason for this is likely that the sound crossing time for gas at scale $k$, $2\pi/k/c_{s}$, decreases towards higher $k$, allowing more time for turbulence to form at smaller scales just after reionization, when the driving force is strongest. This steepens the $k$ dependence of $N_{\tau}$ because smaller scales become turbulent earlier than larger ones.
Figure 9: Estimate of the number of eddy turnover timescales elapsed by $z=4$ as a function of scale. Top Left: comparison of different $z_{\rm re}$ and resolutions. The integrated $N_{\tau}$ is similar in the $z_{\rm re}=7$ and $9$ cases, while in the $z_{\rm re}=5$ case relaxation is still likely ongoing and driving turbulence. We find $N_{\tau}\propto k$ at scales where turbulence is observed in Figure 3, and the high-resolution run traces this scaling up to higher $k$ than the fiducial case. Top Right: dependence on $T_{\min}$. There is little difference with the adiabatic case for $T_{\min}=10^{1}$K, but $N_{\tau}$ decreases and deviates from the expected scaling at higher $T_{\min}$. Bottom Left: dependence on $m_{\rm X}$. All the WDM cases fall well below the CDM case and do not have the expected scaling for turbulence, consistent with Figure 8. Bottom Right: dependence on $v_{\rm bc}^{\rm stream}$. We see much milder dependence here than in the other panels, with $N_{\tau}(k)$ flattening slightly at high $k$ for large $v_{\rm bc}^{\rm stream}$.
We see that our $z_{\rm re}=9$ and $7$ runs yield very similar values of $N_{\tau}$ despite reionizing at different times. This is probably because the turbulence only lasts as long as the driving force is in effect, and this timescale does not vary strongly with redshift. The similarity of the $z_{\rm re}=9$ and $7$ cases suggests that $N_{\tau}$ may be roughly independent of $z_{\rm re}$ when integrated to low enough redshift. In the $z_{\rm re}=5$ case, the driving force probably has not dissipated yet, such that we would find a larger $N_{\tau}$ if we could integrate to lower redshift. For the $z_{\rm re}=7$ and $9$ cases, we find that $N_{\tau}\approx 1$ at our fiducial resolution and $k\approx 800$ hhMpc-1, the highest $k$ for which $N_{\tau}$ traces the gray line. The comparison to our high-resolution run shows that $N_{\tau}$ continues to trace this line up to higher scales when resolution increases, which is encouraging for the prospect of B-field amplification, as we argued in §IV.
The top right panel shows the effect of varying $T_{\min}$. As expected from Figure 8, we find little difference between the adiabatic run and $T_{\min}=10$K. The $T_{\min}=10^{2}$K case deviates from the $\propto k$ appreciably, and the $T_{\min}=10^{3}$K case falls well below $N_{\tau}=1$ at all scales. This confirms our earlier finding that turbulence is suppressed at $T_{\min}=10^{2}$K and disappears for pre-heating rates much higher than this. We see something similar to the $T_{\min}=10^{3}$K case in the lower left panel, which shows results for different WDM particle masses. None of the curves reach $N_{\tau}=1$ or match the scaling seen in the CDM case. The lower right panel shows that $v_{\rm bc}^{\rm stream}$ has a relatively small effect on the inferred $N_{\rm tau}$, but does modestly change its scaling with $k$.
C.2 Sensitivity to the self-shielding treatment
We test here whether our findings are sensitive to the implementation of self-shielding in our simulations. Throughout this work, we use the same setup for sending I-fronts across RT domains described in Cain et al. [17]. At $z_{\rm re}$, we begin tracing rays from the boundaries of RT domains, and “freeze” both cosmic time and hydrodynamical evolution until the I-fronts have crossed the RT domains. This ensures that the gas throughout the box responds coherently to the effects of reionization, and that the volume is fully reionized at exactly $z_{\rm re}$. During this time, the reduced speed of light is set to $5\times$ the speed of the I-front crossing the RT domain, or its full value, whichever is smaller, to ensure the correct I-front speed and post-I-front gas temperature [2, 3]. At $z<z_{\rm re}$, we use a reduced speed of light of $\tilde{c}=2\times 10^{-3}$, which we find does not affect the self-shielding properties of the gas significantly [^6] Because of the tiny light-crossing time at the cell size, these simulations become computationally intractable if the speed of light is not reduced significantly. . We show here that this procedure does not affect our qualitative results, but does simplify their interpretation.
We have run two additional simulations with $N=1024^{3}$. The first flash-ionizes the box at $z=z_{\rm re}$ (just like our $N=2048^{3}$) run, and thus has no self-shielding. The second sends a single plane-parallel I-front across the entire box at $z=z_{\rm re}$ from just one direction, and does not freeze the density field during I-front passage. Respectively, these two approaches minimize and maximize the effect of self-shielding on IGM gas dynamics.
C.3 Test with Nyx
Figure 11: Comparison of IGM turbulence in RadHydro and Nyx. We show slices through the temperature map at $z=6$, 55, and 44 (left to right) for our fiducial RadHydro run, a Nyx run with matching resolution, and our highest resolution RadHydro run (top to bottom). We find qualitatively similar turbulence in Nyx to that observed in RadHydro. Indeed, Nyx shows more turbulence on smaller scales at a fixed resolution than RadHydro, suggesting its algorithm is less diffusive.